EAC

You are evaluating two different silicon wafer milling machines. The Techron I costs $249,000, has a three-year life, and has pretax operating costs of $66,000 per year. The Techron II costs $435,000, has a five-year life, and has pretax operating costs of $39,000 per year. For both milling machines, use straight-line depreciation to zero over the project’s life and assume a salvage value of $43,000. If your tax rate is 22 percent and your discount rate is 11 percent, compute the EAC for both machines


Techron I=

Techron II=
Answer #1

The EAC of Techron I can be calculated using the following inputs on the spreadsheet:

1 Techron 249000 Techron 11 435000 =B2/B3 43000 66000 0.1 0.35 =C2/C3 43000 39000 2 Cost 3 Life 4 Calculate depreciation as fThe results obtained are as follows:

DE 1 B C Techroni Techron II $249,000.00 $435,000.00 $83,000.00 $43,000.00 $66,000.00 10% 35% $87,000.00 $43,000.00 $39,000.0Thus, the EAC of techron I is -$100,985.65.

Calculate the EAC of Techron II as follows:

D E F G Techroni 249000 Techron II 435000 5 =B2/B3 43000 66000 0.1 10.35 =C2/C3 43000 39000 1 2 Cost 3 Life 4. Calculate deprThe results obtained are as follows:

B E F G Techroni Techron II $249,000.00 $435,000.00 $83,000.00 $43,000.00 $66,000.00 10% 35% $87,000.00 $43,000.00 $39,000.00

Thus, the EAC of techron II is -$102,608.61.